OpenStudy (anonymous):

If the probability that a certain tennis player will serve an ace is 1/10, what is the probability that he will serve exactly 2 aces out of 8 serves? (Round your answer to four decimal places.)

OpenStudy (kropot72):

This is solved using the binomial distribution. Do you know the binomial distribution?

OpenStudy (kropot72):

@anc36 Are you there?

OpenStudy (anonymous):

no, what's that?

OpenStudy (kropot72):

\[\left(\begin{matrix}n \\ r\end{matrix}\right)p ^{}(1-p)^{n-r}\] where p = probability of success = 0.1 n = number of trials = 8 r = number of sucesses = 2 Substituting these values in the expression above gives \[P(2\ aces)=(\frac{8\times 7}{2})\times (0.1)^{2}\times (0.9)^{6}=You\ can\ calculate\]